For the following reaction, C2H4(g) + H2(g) ↔ C2H6(g) determine the value of ∆rG (in kJ mol-1) at 20 °C given that ∆rG° = -92.2 kJ mol-1, P(C2H4) = 1.41×10-2 bar, P(H2) = 1.47×10-2 bar, and P(C2H6) = 3.00×10-1 bar.
Answer #1

..shwat CH4 (9) + H₂ (9) = CHb (g) & PCH 3x100 (Ena) (Pre) (1.41x10²) (1.47*18²) 8 = 1447.4 AG = AG + RT lua = -92.2 x103 + 8